hello Please Solve this question


Two identical particles,each having a charge of 2.0 X 10-4 C and mass of 10 g,are kept at a seperation of 10cm and then released.What would bee the speeds of the particle When the seperation becomes large ?


Answer :600m/s

Grade:12

1 Answers

vikas askiitian expert
13 years ago

   

 

 

initial potential energy is (PE)i = kq2 /r                  (charges are of same magnitude q)

                    k = 9 . 109 units

                                        (PE)i  =3600j                 (after substituting value of charge and K)

finally when distance is maximum then (PE)=kq2 /r                    

                                r=r(max)=infinity

      so final potential energy is 0...

initially kinetic energy is zero ....

finally kinetic energy is (KE)f =  mv12 /2   +  mv2 2 /2

                                         =m(v12 + v22 )/2 ................1

from conservation of energy ,total energy initial = total energy final

    therefore ,     m(v12 +v22 )/2 =3600 ..................2

       since no external; force is acting so momentam will be conserved..

     initial momentam =0

    final momentam=mv1 +mv2

      mv1 +mv2 =0

            v1=-v2 ..................1

    now from eq 2 putting v1=-v2

     v1=600m/s

      v2=-600m/s